# Frenet Immersed Finite Element Spaces on Triangular Meshes

### v1

@@ Abstract.
In this paper, we develop geometry-conforming immersed finite element (GC-IFE) spaces on triangular meshes for elliptic interface problems. These IFE spaces are constructed via a Frenet–Serret mapping that transforms the interface curve into a straight line, allowing the interface jump conditions to be imposed exactly. Extending the framework of [2024AdjeridLinMeghaichi, 2025AdjeridLinMeghaichi2] from rectangular to triangular meshes, we introduce three procedures for constructing high-degree Frenet-IFE spaces: an initial construction based on monomial bases, a generalized construction using orthogonal polynomials, and reconstruction methods aimed at improving the conditioning of the associated mass matrix. The optimal approximation capability of the proposed IFE spaces is demonstrated through numerical examples. We further incorporate these spaces into interior penalty discontinuous Galerkin methods for elliptic interface problems and observe optimal convergence rates in the $H^{1}$ and $L^{2}$ norms.
@@ 1. Introduction
In this paper, we develop a high-degree geometry-conforming immersed finite element (GC-IFE) on triangular meshes for the second-order elliptic interface problem:
@@
where $\left\llbracket v\right\rrbracket_{\Gamma}:=v^{+}|_{\Gamma}-v^{-}|_{\Gamma}$ is the jump across the interface, and $\mathbf{n}$ denotes the normal on $\Gamma$ from $\Omega^{-}$ to $\Omega^{+}$. As in [2024AdjeridLinMeghaichi, 2017AdjeridGuoLin, 2019GuoLin2], for the construction of an IFE space with a high-degree polynomial space $\mathbb{P}_{m}$, we further assume that the solution $u$ satisfies the Laplacian extended jump conditions:
@@
The interface problem (1.1)- (1.5) arises widely in science and engineering applications. Among related numerical methods, the Immersed Finite Element (IFE) method [ZLi_TLin_XWu, 2004LiLinLinRogers, 2015LinLinZhang, 2019LinSheenZhang, 2005KafafyLinLinWang, 2021GuoZhang] is a class of numerical techniques designed to solve such problems without requiring the mesh to align with the interface. Its key idea is to modify the local shape functions only on interface elements according to the prescribed jump conditions, while leaving those on non-interface elements unchanged.
@@
Traditionally, local IFE functions are constructed as piecewise polynomials on subelements separated by the interface. For low-order IFE spaces [2004LiLinLinRogers, 2008HeLinLin, 2019GuoLin], the interface can be adequately represented by a linear approximation, and the associated IFE spaces are built accordingly. However, for high-degree IFE spaces, such a linear approximation becomes insufficient. When the interface is an arbitrary curve, the jump conditions cannot be enforced exactly because polynomials pieces cannot match perfectly along the interface curve. As a result, the jump conditions are often imposed on the interface curve only in a weak sense. Examples include least-squares enforcement of the jump conditions [2017AdjeridGuoLin, 2021ChenZhang, 2025ChenZhang], Cauchy extension [2019GuoLin2, 2023AdjeridBabuskaGuoLin], and enforcing jump conditions on selected quadrature points [2022Adjerid]. While these approaches yield IFE spaces with sufficient approximation capability, the resulting IFE spaces are geometrically nonconforming, meaning the local IFE functions generally are not in $H^{1}$. This drawback has two main ramifications. First, an appropriate penalty must be incorporated along the interface to mitigate the effects of the inherent discontinuity of the IFE functions [2023AdjeridBabuskaGuoLin, 2026ZhuangZhangSarkisLin], which consequently increases the complexity of the IFE scheme. Second, the coexistence of discontinuity and penalty terms along the interface poses additional challenges for the associated error analysis.
@@
Recently, geometrically conforming IFE spaces were introduced on rectangular meshes [2024AdjeridLinMeghaichi]. These IFE space are constructed with the Frenet-Serret transformation, an idea from differential geometry, to map an arbitrary interface curve in the Cartesian $x$-$y$ plane into a straight line segment in the Frenet coordinates $\eta$-$\xi$. Although standard tensor-product polynomial spaces $\mathbb{Q}_{m}$ are used on Frenet reference elements, the corresponding IFE shape functions on the physical element are no longer piecewise polynomials because of the nonlinearity of the Frenet mapping. Nevertheless, these new IFE functions preserve continuity inside each interface element, which makes the resulting space geometry-conforming. In [2025AdjeridLinMeghaichi], the approximation properties of the Frenet-IFE spaces were theoretically established and the optimal a priori error estimates for associated immersed discontinuous Galerkin solution are proved. In [2025AdjeridLinMeghaichi2], reconstruction techniques for Frenet-IFE bases were further developed to improve the conditioning of the associated mass matrix.
@@
The Frenet-IFE framework [2024AdjeridLinMeghaichi, 2025AdjeridLinMeghaichi, 2025AdjeridLinMeghaichi2] has so far been developed exclusively for domains that can be discretized by rectangular meshes. This restriction limits its applicability, since many practical problems involve domains with complex geometries. In particular, rectangular meshes cannot be constructed on non-rectangular polygonal domains or domain with curved boundaries. In contrast, triangular meshes provide much greater flexibility: they can naturally conform to complex boundaries, and enable efficient local mesh refinement. These advantages motivate us to generalize and extend the Frenet-IFE framework from Cartesian rectangular meshes to unstructured triangular meshes.
@@
This paper is strongly influenced by, and builds upon, the developments in [2025AdjeridLinMeghaichi2]. We present three approaches for constructing high-degree, geometry-conforming IFE spaces on triangular meshes. The first approach, referred to as initial construction, uses $(m+1)(m+2)/2$ monomials forming the basis of $\mathbb{P}_{m}$. Thanks to its simple structure, $m(m+1)/2$ of the resulting IFE basis functions can be written down explicitly, while the remaining $m+1$ functions are determined by solving a small linear system. The second approach, called the general construction, uses orthogonal polynomials to build IFE basis functions, not only providing a more general framework but also facilitating the assembly of local matrices. The third approach, referred to as reconstruction, employs a singular value decomposition to produce an enhanced Frenet-IFE basis that yields optimally conditioned mass matrices, which is essential for numerical stability in solving time-dependent interface problems.
@@
The remainder of the paper is organized as follows. Section 2 reviews the necessary notations, the Frenet transformation, and related results previously established for rectangular meshes. Sections 3 and 4 present the initial and general constructions of the Frenet-IFE basis functions on triangular meshes. In Section 5, we apply the reconstruction strategies introduced in [2025AdjeridLinMeghaichi2] to IFE basis functions on triangular mesh for improving the conditioning of mass matrices. In Section 6, we report numerical results for the $L^{2}$ projection and discontinuous Galerkin approximation of the elliptic interface problems. Finally, Section 7 provides some concluding remarks.
@@ 2. Preliminaries
In this section, we review Frenet transformation and recall some related notations and formula previously presented in [2024AdjeridLinMeghaichi, 2025AdjeridLinMeghaichi, 2025AdjeridLinMeghaichi2].
Let $\mathcal{T}_{h}=\{K\}$ be a shape-regular triangulation of $\Omega$ that is independent of the interface $\Gamma$. We call $K$ an interface element if $\mathring{K}\cap\Gamma=\emptyset$, and denote by $\mathcal{T}_{h}^{i}$ the collection of all interface elements. The set of noninterface elements is then $\mathcal{T}_{h}^{n}=\mathcal{T}_{h}/\mathcal{T}_{h}^{i}$.
@@
In differential geometry, the Frenet–Serret apparatus [gray_2006, ONeill_DiffGeom_2010] associated with the curve $\mathbf{g}(\xi)$ consists of the unit tangent vector $\bm{\tau}(\xi)$, the unit normal vector $\bm{n}(\xi)$, and the curvature $\kappa(\xi)$ defined by
$\bm{\tau}(\xi)=\frac{1}{\|\mathbf{g}^{\prime}(\xi)\|}\mathbf{g}^{\prime}(\xi),\qquad\bm{n}(\xi)=Q\bm{\tau}(\xi),\qquad\kappa(\xi)=\frac{1}{\|\mathbf{g}^{\prime}(\xi)\|^{3}}\big(\mathbf{g}^{\prime}(\xi)^{T}Q\mathbf{g}^{\prime\prime}(\xi)\big)$ (2.7)
@@
The Frenet transformation $P_{\Gamma}:(\eta,\xi)\to(x,y)$ is defined by
@@
within which the transformation $P_{\Gamma}$ is a bijection. Hence its inverse
@@
For sufficiently small mesh size $h$, every interface element in $\mathcal{T}_{h}$ lies inside $N_{\Gamma}(\epsilon)$. Consider an interface element $K\in\mathcal{T}_{h}^{i}$ with vertices $A_{1}$, $A_{2}$, and $A_{3}$. Let $R_{\Gamma}(A_{i})=[\eta_{i},\xi_{i}]$. We form a fictitious element $K_{F}$ containing $K$, with two curved edges parallel to the interface $\Gamma$ and two straight edges. By the inverse transformation, the interface triangle $K$ becomes a curved triangle $\hat{K}=R_{\Gamma}(K)\subset\widehat{K}_{F}=R_{\Gamma}(K_{F})$. A key feature of the Frenet transformation is that the interface segment $\Gamma_{K_{F}}=\Gamma\cap K_{F}$ is mapped to the vertical line segment $\widehat{\Gamma}_{K_{F}}=R_{\Gamma}(\Gamma_{K_{F}})$ in the $(\eta,\xi)$-plane. See Figure 1 for an illustration.
@@
With ${\hat{\beta}}=\beta\circ P_{\Gamma}$ and using this Frenet transformation, the interface jump conditions in (1.3)-(1.5) become
@@
where the Laplacian expressed in Frenet coordinates is given by
@@
$J_{0}(\eta,\xi)=\rho^{2}(\eta,\xi),\quad J_{1}(\eta,\xi)=\kappa(\xi)\psi(\eta,\xi),\quad J_{2}(\eta,\xi)=\rho^{2}(\eta,\xi)\left(\eta\kappa^{\prime}(\xi)\psi(\eta,\xi)\frac{\mathbf{g}^{\prime}(\xi)\cdot\mathbf{g}^{\prime\prime}(\xi)}{\|\mathbf{g}^{\prime}(\xi)\|^{2}}\right).$ (2.13)
Following the ideas in [2018AdjeridBenromdhaneLin], the extended jump conditions (2.11) lead to the following weak jump conditions on ${\hat{\Gamma}}_{K_{F}}$:
$\int_{{\hat{\Gamma}}_{K_{F}}}{\hat{v}}\left\llbracket{\hat{\beta}}\partial_{\eta}^{j}\mathscr{L}({\hat{u}})\right\rrbracket_{{\hat{\Gamma}}_{K_{F}}}d\xi$ $=$ $0,\quad\forall{\hat{v}}\in\mathbb{P}_{m-2-j}({\hat{\Gamma}}_{K_{F}}),~~~~~~j=0,1,\cdots,m-2.$ (2.14)
On each Frenet reference element $\hat{K}_{F}$, we will construct the piecewise polynomial space
@@
where $\mathbb{P}_{m}$ is the polynomial space of degree at most $m$. The local geometry-conforming IFE (GC-IFE) space on the original interface triangle $K$ is defined by
${\mathcal{V}}^{m}_{\beta}({K})=\left\{\phi=\hat{\phi}\circ R_{\Gamma}|_{K}:~\hat{\phi}\in\hat{\mathcal{V}}^{m}_{\hat{\beta}}(\hat{K}_{F})\right\}.$ (2.16)
Then the global GC-IFE space based on the polynomial space $\mathbb{P}_{m}$ is defined as
@@
In the following three sections, we will present procedures to construct the GC-IFE space ${\mathcal{V}}^{m}_{\beta}({K})$ on each interface element $K\in\mathcal{T}_{h}$.
@@ 3. Initial Construction of IFE Spaces on Triangles
The first procedure for constructing the $\mathbb{P}_{m}$ Frenet-IFE space on a triangular interface element follows the initial construction procedure of the $\mathbb{Q}_{m}$ Frenet-IFE space on a rectangular interface element in [2024AdjeridLinMeghaichi]. We begin by constructing an IFE basis in Frenet coordinates $\eta$-$\xi$ on ${\hat{K}}_{F}$. Then, through a change of variable with the Frenet transformation, we obtain a set of IFE basis functions on fictitious element $K_{F}$. Finally, the Frenet-IFE functions based on $\mathbb{P}_{m}$ polynomials are obtained by restricting these functions on $K_{F}$ to the triangular interface element $K$.
Let $\left\{p_{i}(\xi)\right\}_{i=0}^{m}$ be a basis of $\mathbb{P}_{m}({\hat{\Gamma}}_{K_{F}})$. It is straightforward to verify that the following polynomials form a basis for $\mathbb{P}_{m}({\hat{K}}_{F})$:
$\begin{array}[]{ccccccccc}p_{0}(\xi)&\\ p_{1}(\xi)&\eta p_{0}(\xi)&\\ p_{2}(\xi)&\eta p_{1}(\xi)&\eta^{2}p_{0}(\xi)\\ p_{3}(\xi)&\eta p_{2}(\xi)&\eta^{2}p_{1}(\xi)&\eta^{3}p_{0}(\xi)\\ \vdots&\vdots&\vdots&\vdots&\ddots\\ p_{m}(\xi)&\eta p_{m-1}(\xi)&\eta^{2}p_{m-2}(\xi)&\eta^{3}p_{m-3}(\xi)&\cdots&\eta^{m}p_{0}(\xi)\end{array}$ (3.24)
For $m\geq 2$, there are $l_{m}={(m-1)m}/{2}$ polynomials listed between the third and $(m+1)$-th columns in the table (3.24). For convenience, we denote these basis polynomials by $\mathcal{N}_{i}(\eta,\xi)$ with $i=1,2,\cdots,l_{m}$ where
@@
Following the same argument used in the proof of Lemma 3 of [2024AdjeridLinMeghaichi] we obtain the analogous result for $\mathbb{P}_{m}$ polynomial spaces.
Lemma 3.1.
@@
Using the basis in (3.24), we consider two polynomials ${\hat{\phi}}^{s}(\eta,\xi)\in\mathbb{P}_{m}({\hat{K}}_{F}^{s}),~s=\pm$ expressed as follows:
@@
In the discussions from now on, for each interface element $K$, we let $K^{s}=K\cap\overline{\Omega^{s}},~s=\pm$. Also, for $s=\pm$, we have $s^{\prime}=\mp$. Following the idea of Lemma 1 in [2024AdjeridLinMeghaichi], we obtain the unisolvency property.
Lemma 3.2.
Given a polynomial ${\hat{\phi}}^{s}(\eta,\xi)\in\mathbb{P}_{m}({\hat{K}}_{F}^{s})$, there exists a unique ${\hat{\phi}}^{s^{\prime}}(\eta,\xi)\in\mathbb{P}_{m}({\hat{K}}_{F}^{s^{\prime}})$ such that
@@
${\hat{\phi}}_{0,i}(\eta,\xi)=\begin{cases}{\hat{\phi}}_{0,i}^{-}(\eta,\xi)=p_{i}(\xi),&(\eta,\xi)\in{\hat{K}}_{K}^{-},\\[5.0pt] {\hat{\phi}}_{0,i}^{+}(\eta,\xi)=\sum\limits_{k=0}^{m}\eta^{k}\left(\sum\limits_{l=0}^{m-k}C_{k,l}^{+}p_{l}(\xi)\right),&(\eta,\xi)\in{\hat{K}}_{K}^{+},\end{cases}$ (3.27)
where coefficients $C_{k,l}^{+}$, $0\leq l\leq m-k,~0\leq k\leq m$ are to be determined so that ${\hat{\phi}}_{0,i}(\eta,\xi)$ satisfies the interface jump conditions (2.9), (2.10), and (2.14). The jump condition (2.9) implies that
$p_{i}(\xi)={\hat{\phi}}_{0,i}^{-}(0,\xi)={\hat{\phi}}_{0,i}^{+}(0,\xi)=\sum\limits_{l=0}^{m}C_{0,l}^{+}p_{l}(\xi),~\forall\xi\in{\hat{\Gamma}}_{K_{F}}.$
@@
${\hat{\phi}}_{0,i}^{+}(\eta,\xi)=p_{i}(\xi)+\sum\limits_{k=1}^{m}\eta^{k}\left(\sum\limits_{l=0}^{m-k}C_{k,l}^{+}p_{l}(\xi)\right),\qquad\frac{\partial{\hat{\phi}}_{0,i}^{+}(\eta,\xi)}{\partial\eta}=\sum\limits_{k=1}^{m}k\eta^{k-1}\left(\sum\limits_{l=0}^{m-k}C_{k,l}^{+}p_{l}(\xi)\right).$

### v2

@@ Abstract.
In this paper, we develop geometry-conforming immersed finite element (IFE) spaces on triangular meshes for elliptic interface problems. The construction is built on a Frenet–Serret mapping that transforms a smooth interface curve into a straight line, so that the interface jump conditions can be imposed exactly. Extending the framework of [2025AdjeridLinMeghaichi] from rectangular meshes to triangular meshes, we introduce three types of high-order Frenet-IFE constructions: an initial construction using monomial bases, a general construction using orthogonal polynomials, and reconstructed IFE bases designed to improve the conditioning of the mass matrix. The approximation properties of these new IFE spaces are investigated through extensive numerical experiments. We also incorporate the new IFE spaces into interior penalty discontinuous Galerkin methods for solving elliptic interface problems, and demonstrate optimal convergence rates in $H^{1}$- and $L^{2}$- norms.
Key words and phrases:
Interface problems, Immersed finite element, Frenet mapping, Triangular mesh, Conditioning.
2000 Mathematics Subject Classification:
35R05, 65N15, 65N30
@@ 1. Introduction
In this paper, we develop a high-order geometry-conforming immersed finite element (GC-IFE) on triangular meshes for the second-order elliptic interface problem:
@@
where $\left\llbracket v\right\rrbracket_{\Gamma}:=v^{+}|_{\Gamma}-v^{-}|_{\Gamma}$ is the jump across the interface, and $\mathbf{n}$ denotes the unit normal on $\Gamma$ from $\Omega^{-}$ to $\Omega^{+}$. As described in [2017AdjeridGuoLin, 2024AdjeridLinMeghaichi, 2019GuoLin2], with the intention of constructing a high-order $\mathbb{P}_{m}$ IFE space, we further assume that the solution $u$ satisfies the Laplacian extended jump conditions:
@@
The interface problem (1.1) - (1.5) arises widely in science and engineering applications. The Immersed Finite Element (IFE) method [2019GuoLinZhuang, 2021GuoZhang, 2026JiLi, 2023JiWangChenLi, 2005KafafyLinLinWang, 1998Li, 2015LinLinZhang, 2019LinSheenZhang] is a class of numerical techniques designed to solve such problems without requiring the mesh to align with the interface. Its key idea is to modify the local shape functions only on interface elements according to the prescribed jump conditions, while leaving the finite element space unchanged on regular elements.
@@
Traditionally, local IFE functions are constructed as piecewise polynomials on subelements separated by the interface. For low-order IFE spaces [2019GuoLin, 2008HeLinLin, 2022JiWangChenLi2, 2004LiLinLinRogers], the interface can be adequately represented by a linear approximation, and the associated IFE spaces are built accordingly. However, for high-order IFE spaces, such a linear approximation becomes insufficient. When the interface is an arbitrary smooth curve, the jump conditions cannot be enforced exactly because polynomial pieces cannot match perfectly along the interface curve. As a result, the jump conditions are often imposed on the interface curve only in a weak sense. Examples include least-squares enforcement of the jump conditions [2017AdjeridGuoLin, 2021ChenZhang, 2024ChenZhang, 2025ChenZhang], Cauchy extension [2023AdjeridBabuskaGuoLin, 2019GuoLin2], and enforcement of jump conditions at selected points [2022Adjerid, 2014AdjeridBenromdhaneLin, 2017GuzmanSanchezSarkis]. While these approaches yield IFE spaces with sufficient approximation capability, the resulting spaces are geometrically nonconforming, meaning the local IFE functions generally do not lie in $H^{1}$. This drawback has two main ramifications. First, an appropriate penalty must be incorporated along the interface to mitigate the effect of the inherent discontinuity of the IFE functions [2023AdjeridBabuskaGuoLin, 2026ZhuangZhangSarkisLin], which consequently increases the complexity of the IFE scheme. Second, the coexistence of discontinuity and penalty terms along the interface poses additional challenges for the associated error analysis.
@@
Recently, a geometrically conforming IFE space was introduced on rectangular meshes [2024AdjeridLinMeghaichi]. This construction uses the Frenet-Serret transformation, an idea from differential geometry, to map an arbitrary interface curve in the Cartesian $x$-$y$ plane into a straight line segment in the Frenet coordinates $\eta$-$\xi$. Although standard tensor-product polynomial spaces $\mathbb{Q}_{m}$ are used on Frenet reference elements, the corresponding IFE shape functions on the physical element are no longer piecewise polynomials because of the nonlinearity of the Frenet mapping. Nevertheless, these new IFE functions preserve continuity inside each interface element, which makes the resulting space geometry-conforming. In [2025AdjeridLinMeghaichi], the approximation properties of the IFE spaces were theoretically established and the optimal a priori error estimates for immersed discontinuous Galerkin solution were proved. In [2025AdjeridLinMeghaichi2], some reconstruction techniques for Frenet-IFE bases were further developed to improve the conditioning of the associated mass matrix.
@@
The Frenet-IFE framework [2024AdjeridLinMeghaichi, 2025AdjeridLinMeghaichi2, 2025AdjeridLinMeghaichi] has so far been developed exclusively for structured domains discretized by rectangular meshes. This restriction limits its applicability, since many practical problems involve domains with complex geometries. In particular, rectangular meshes cannot be constructed on non-rectangular polygonal domains or domains with curved boundaries. In contrast, triangular meshes provide much greater flexibility. They can naturally conform to complex boundaries, and enable efficient local mesh refinement. These advantages motivate us to generalize and extend the Frenet-IFE framework from Cartesian rectangular meshes to unstructured triangular meshes.
@@
This paper is strongly influenced by, and builds upon, the developments in [2025AdjeridLinMeghaichi2]. We present three approaches for constructing high-order, geometrically conforming IFE spaces on triangular meshes. The first approach, referred to as initial construction, uses $(m+1)(m+2)/2$ monomials forming the basis of $\mathbb{P}_{m}$. Thanks to its simple structure, $m(m+1)/2$ of the resulting IFE basis functions can be written down explicitly, while the remaining $m+1$ functions are determined by solving a small linear system. The second approach, called general construction, uses orthogonal polynomials to build IFE basis functions, providing a more flexible framework. The third approach, named reconstruction, modifies the Frenet IFE basis to achieve optimally conditioned mass matrices, which is crucial for numerical stability, especially for time dependent problems.
@@
The remainder of the paper is organized as follows. Section 2 reviews the necessary notations, the Frenet transformation, and related results previously established for rectangular meshes. Sections 3 and 4 present the initial and general constructions of the Frenet-IFE basis functions on triangular meshes. In Section 5, we apply the reconstruction strategies [2025AdjeridLinMeghaichi2] to IFE basis functions on triangular meshes for improving the conditioning of mass matrices. In Section 6, we report numerical results for the $L^{2}$ projection and discontinuous Galerkin approximation of the elliptic interface problem. Finally, Section 7 provides some concluding remarks.
@@ 2. Preliminaries
In this section, we review some notation for Frenet transformation and recall related results for the rectangular meshes previously presented in [2024AdjeridLinMeghaichi, 2025AdjeridLinMeghaichi2, 2025AdjeridLinMeghaichi].
Let $\mathcal{T}_{h}=\{K\}$ be a shape-regular triangulation of $\Omega$ that is independent of the interface $\Gamma$. We call $K$ an interface element if $K\cap\Gamma{\color[rgb]{0,0,0}\neq}\emptyset$, and denote by $\mathcal{T}_{h}^{i}$ the collection of all interface elements. The set of noninterface elements is then $\mathcal{T}_{h}^{n}=\mathcal{T}_{h}\setminus\mathcal{T}_{h}^{i}$.
@@
The Frenet–Serret apparatus associated with $\mathbf{g}(\xi)$ consists of the unit tangent vector $\bm{\tau}(\xi)$, the unit normal vector $\bm{n}(\xi)$, and the curvature $\kappa(\xi)$ defined by
$\bm{\tau}(\xi)=\frac{\mathbf{g}^{\prime}(\xi)}{\|\mathbf{g}^{\prime}(\xi)\|},\quad\bm{n}(\xi)=Q\bm{\tau}(\xi),\quad\kappa(\xi)=\frac{\mathbf{g}^{\prime}(\xi)^{T}Q\mathbf{g}^{\prime\prime}(\xi)}{\|\mathbf{g}^{\prime}(\xi)\|^{3}}$ (2.2)
@@
The mapping $P_{\Gamma}:(\eta,\xi)\to(x,y)$ is defined by
@@
within which the transform $P_{\Gamma}$ is a bijection. Hence its inverse
@@
For sufficiently small mesh size $h$, every interface element in $\mathcal{T}_{h}$ lies inside $N_{\Gamma}(\epsilon)$. Consider an interface element $K\in\mathcal{T}_{h}^{i}$ with vertices $A_{1}$, $A_{2}$, and $A_{3}$, and let $R_{\Gamma}(A_{i})=[\eta_{i},\xi_{i}]$. Let $\xi_{0,K}=\min\limits_{1\leq i\leq 3}\{\xi_{i}\}$ and $\xi_{1,K}=\max\limits_{1\leq i\leq 3}\{\xi_{i}\}$. We form a fictitious element $K_{F}$ containing $K$, with two curved edges parallel to the interface $\Gamma$ and two straight edges. By applying the inverse mapping, the interface triangle $K$ becomes a curved triangle $\hat{K}=R_{\Gamma}(K)\subset\widehat{K}_{F}$. A key feature of the Frenet mapping is that the interface segment $\Gamma_{K_{F}}=\Gamma\cap K_{F}$ is mapped to the vertical line segment $\widehat{\Gamma}_{K_{F}}=R_{\Gamma}(\Gamma_{K_{F}})$ in the $(\eta,\xi)$-plane.
See Figure 1 for an illustration.
@@
Using this coordinate transformation, every function $v:N_{\Gamma}(\epsilon)\rightarrow\mathbb{R}$ is associated with the following function of the Frenet coordinates:
${\hat{v}}=v\circ P_{\Gamma}:[-\epsilon,\epsilon]\times[\xi_{s},\xi_{e}]\to\mathbb{R}.$ (2.4)
Moreover, the interface jump conditions in (1.3)-(1.5) are transformed to
@@
where the Laplacian in Frenet coordinate is given by
@@
$\begin{split}J_{0}(\eta,\xi)&=\rho^{2}(\eta,\xi),\quad J_{1}(\eta,\xi)=\kappa(\xi)\psi(\eta,\xi),\\ J_{2}(\eta,\xi)&=\rho^{2}(\eta,\xi)\left(\eta\kappa^{\prime}(\xi)\psi(\eta,\xi)\frac{\mathbf{g}^{\prime}(\xi)\cdot\mathbf{g}^{\prime\prime}(\xi)}{\|\mathbf{g}^{\prime}(\xi)\|^{2}}\right).\end{split}$ (2.9)
and
$\rho(\eta,\xi)=\|\mathbf{g}^{\prime}(\xi)\|^{-1}\psi(\eta,\xi),\qquad\psi(\eta,\xi)=\frac{1}{1+\eta\kappa(\xi)}.$
Following the ideas in [2018AdjeridBenromdhaneLin], the extended jump conditions (1.5), rewritten in Frenet coordinates as (2.7), lead to the following weak jump conditions on ${\hat{\Gamma}}_{K_{F}}$:
$\int_{{\hat{\Gamma}}_{K_{F}}}{\hat{v}}\left\llbracket{\hat{\beta}}\partial_{\eta}^{j}\mathscr{L}({\hat{u}})\right\rrbracket_{{\hat{\Gamma}}_{K_{F}}}d\xi=0,\quad\forall{\hat{v}}\in\mathbb{P}_{m-2-j}({\hat{\Gamma}}_{K_{F}}),$ (2.10)
for $j=0,1,\cdots,m-2.$
On each Frenet reference element $\hat{K}_{F}$ we construct the piecewise polynomial space
@@
where $\mathbb{P}_{m}$ is the polynomial space of order no more than $m$. The local geometric-conforming IFE space on the original interface triangle $K$ is defined by
${\mathcal{V}}^{m}_{\beta}({K})=\left\{\hat{\phi}\circ R_{\Gamma}|_{K}:~\hat{\phi}\in\hat{\mathcal{V}}^{m}_{\hat{\beta}}(\hat{K}_{F})\right\}.$ (2.12)
Then the global $\mathbb{P}_{m}$ GC-IFE space is defined as
@@
In the following three sections, we introduce several strategies to construct $\mathbb{P}_{m}$ GC-IFE space ${\mathcal{V}}^{m}_{\beta}({K})$ on interface triangles.
@@ 3. Initial Construction of IFE Spaces on Triangles
The first strategy to build $\mathbb{P}_{m}$ Frenet-IFE spaces on triangles follows the initial construction of the $\mathbb{Q}_{m}$ Frenet-IFE spaces on rectangular elements in [2024AdjeridLinMeghaichi]. We begin by constructing a monomial basis in $\mathbb{P}_{m}$ in Frenet coordinates $\eta$-$\xi$ on ${\hat{K}}_{F}$. Then, through a change of variable with the Frenet transform, we obtain a set of IFE basis functions on fictitious element $K_{F}$. Finally, the Frenet IFE functions based on $\mathbb{P}_{m}$ polynomials are obtained by restricting these functions on $K_{F}$ to the triangular interface element $K$.
Let $\left\{p_{i}(\xi)\right\}_{i=0}^{m}$ be a basis of $\mathbb{P}_{m}(\widehat{\Gamma}_{K_{F}})$ with $\text{deg}(p_{i})=i,~0\leq i\leq m$. It is straightforward to verify that the following polynomials form a basis for $\mathbb{P}_{m}({\hat{K}}_{F})$:
$\begin{array}[]{ccccccccc}p_{0}(\xi)&\\ p_{1}(\xi)&\eta p_{0}(\xi)&\\ p_{2}(\xi)&\eta p_{1}(\xi)&\eta^{2}p_{0}(\xi)\\ p_{3}(\xi)&\eta p_{2}(\xi)&\eta^{2}p_{1}(\xi)&\eta^{3}p_{0}(\xi)\\ \vdots&\vdots&\vdots&\vdots&\ddots\\ p_{m}(\xi)&\eta p_{m-1}(\xi)&\eta^{2}p_{m-2}(\xi)&\eta^{3}p_{m-3}(\xi)&\cdots&\eta^{m}p_{0}(\xi).\end{array}$ (3.7)
There are $l_{m}={(m-1)m}/{2}$ polynomials in columns $3$ through $(m+1)$ of the table (3.7). For convenience, we denote these basis polynomials by $\mathcal{N}_{i}(\eta,\xi)$ with $i=1,2,\cdots,l_{m}$ where
@@
Following the same argument used in proving Lemma 3 of [2024AdjeridLinMeghaichi] we obtain the analogous result for $\mathbb{P}_{m}$ polynomial spaces.
Lemma 3.1
@@
Using the basis in (3.7), we can expand ${\hat{\phi}}^{s}(\eta,\xi)\in\mathbb{P}_{m}({\hat{K}}_{F}^{s}),~s=\pm$ as follows:
@@
Following the idea of Lemma 1 in [2024AdjeridLinMeghaichi], we obtain the unisolvency property.
Lemma 3.2
Given a polynomial ${\hat{\phi}}^{s}(\eta,\xi)\in\mathbb{P}_{m}({\hat{K}}_{F}^{s})$, there exists a unique ${\hat{\phi}}^{s^{\prime}}(\eta,\xi)\in\mathbb{P}_{m}({\hat{K}}_{F}^{s^{\prime}})$ such that the piecewise polynomial $\hat{\phi}$ defined by
@@
${\hat{\phi}}_{0,i}(\eta,\xi)=\begin{cases}{\hat{\phi}}_{0,i}^{-}(\eta,\xi)=p_{i}(\xi),&(\eta,\xi)\in{\hat{K}}_{K}^{-},\\[5.0pt] {\hat{\phi}}_{0,i}^{+}(\eta,\xi)=\sum\limits_{k=0}^{m}\eta^{k}\left(\sum\limits_{l=0}^{m-k}C_{k,l}^{+}p_{l}(\xi)\right),&(\eta,\xi)\in{\hat{K}}_{K}^{+}.\end{cases}$ (3.10)
where coefficients $C_{k,l}^{+}$, $0\leq l\leq m-k,~0\leq k\leq m$ are chosen so that ${\hat{\phi}}_{0,i}(\eta,\xi)$ satisfies the interface jump conditions (2.5), (2.6), and (2.10). The jump condition (2.5) implies that
$p_{i}(\xi)={\hat{\phi}}_{0,i}^{-}(0,\xi)={\hat{\phi}}_{0,i}^{+}(0,\xi)=\sum\limits_{l=0}^{m}C_{0,l}^{+}p_{l}(\xi),~\forall\xi\in\widehat{\Gamma}_{K_{F}}.$
@@
${\hat{\phi}}_{0,i}^{+}(\eta,\xi)$ $=p_{i}(\xi)+\sum_{k=1}^{m}\eta^{k}\left(\sum_{l=0}^{m-k}C_{k,l}^{+}p_{l}(\xi)\right),$
$\frac{\partial{\hat{\phi}}_{0,i}^{+}(\eta,\xi)}{\partial\eta}$ $=\sum_{k=1}^{m}k\eta^{k-1}\left(\sum_{l=0}^{m-k}C_{k,l}^{+}p_{l}(\xi)\right).$